7.6.27 Solve the given initial value problem using the method of Laplace transforms. z + 6z + 8z = e-bu(t-1); Z(0) = 2, z
Answer #1

2+62+82 = é uctt), 2(0) = 2, 2/(0) = -6 Apply leplace transform L2 Z+62+82}= debt uct-1)} +8242} = hebt uiters 222 +62hz25+6 & elst (5+65+8) 2(3) = S46 25+6 + 215) = 6-1816) 5+6 8765+ 8 e-1516) 25+ 6 215) S76518 (5+6) (5+68+8) Apply invers lapte( )، پس = (ft (، ) 4 ( ، ) ، f(t) = 4 fit-a) = f(t-1) - 40 feet in the cerere) - ucev) [fe2c61f2461) ( - - uit ) e-let- 1o-4(